OpenStudy (anonymous):

A car traveled 16 km north and then 20 km east. The total time of travel was 0.75 h. Find the average speed and average velocity of the car for this trip.

OpenStudy (anonymous):

This one I guess

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I need solution.

OpenStudy (anonymous):

Many times speed=velocity. But here I think you need two different things. For the speed: He traveled 16+20=36km in 0.75h 36/0.75= 48km/h For the velocity he needs a vector solution.

OpenStudy (anonymous):

So he traveled North 16km and east 20km Draw this and you will see that the actual distance can be found by using the Pythagorean Theorem. 16^2+20^2=d^2 (d is distance) 656=d^2 d=25.61 Average velocity is 25.61/0.75=34.15 km/h

OpenStudy (anonymous):

Thanks again Andras